OpenStudy (anonymous):

Help please!! If a line passes through the point (-1,2) and has a slope m=a/b, for a>0 and b>0, then which of the following point must lie on the graph of this line? A (-1+b, 2-a) B (-1-a,2-b) C (-1+b, 2+a) (D was a wrong answer)

jimthompson5910 (jim_thompson5910):

y = mx + c y = (a/b)*x + c ... plug in m = a/b 2 = (a/b)*(-1) + c ... plug in (x,y) = (-1,2) now we solve for c 2 = (a/b)*(-1) + c 2 = -a/b + c c = 2 + (a/b) ---------------------------------- So the equation is y = (a/b)x + 2 + (a/b) from here, you test each point by plugging it into the equation above. If you get a true equation, then the point lies on the line.

OpenStudy (anonymous):

the answer I got was c and it was right thanks!

jimthompson5910 (jim_thompson5910):

np